A graded index fiber with a parabolic index profile supports the propagation of




742 guided modes. The fiber has a numerical aperture in air of 0.3 and a core diameter




of 70 μm. Determine the wavelength of the light propagating in the fiber.




Further estimate the maximum diameter of the fiber which gives single-mode




operation at the same wavelength.

1
Expert's answer
2021-12-08T10:00:50-0500

N=742

NA = 0.3

d = 70 "\mu"

Number of modes of propagation in graded index fiber is

"N = 2.45(\frac{d \times NA}{λ})^2 \\

742 = \frac{2.45(70 \times 10^{-6} \times 0.3)^2}{λ^2} \\

λ^2 = \frac{2.45(21 \times 10^{-6})^2}{742} \\

λ = \sqrt{1.456 \times 10^{-12}} \\

λ = 1.2067 \times 10^{-6} \; \mu m"

For single mode operation:

"V≤2.405 \\

\frac{2 \pi \times a \times NA}{λ } ≤ 2.405 \\

a ≤ \frac{2.405 \times λ }{2 \pi \times NA} \\

a ≤ \frac{2.405 \times 1.2067 \times 10^{-6}}{2 \pi \times 0.3} \\

a ≤ 1.54 \times 10^{-6} \; m"

a-radius ≤ 1.54 "\mu" m and diameter d = 2a

"d ≤ 2 \times 1.54 \; \mu m \\

d ≤ 3.08 \; \mu m"

The maximum diameter "d = 3.08 \; \mu m"